# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ III.2.1 Topological amplitudes in $T\to DS$ decays
the mixing angle is taken to be $\phi=174.6^{\circ}$aiani:2004uc . For simplicity of notation, we do not distinguish the topological amplitudes in scenario I and scenario II, which, however, do not imply that they are identical. We follow refs. Jaffe:1976ig ; Jaffe:1976ih on the quark contents of these low-lying scalar particles. For more information on the flavor structure of these particles, see Appendix B.
@@ III.2.2 $T\to DS$ decay amplitudes
$1^{*}$ $T^{++}_{c\bar{s}}\to D^{+}\kappa^{+}$ $2C_{S}+C_{DS}$ $T^{++}_{c\bar{s}}\to D_{s}^{+}a_{0}^{+}$ $-(2C_{S}+C_{DS})$
$2^{*}$ $T^{+}_{c\bar{s}}\to D^{0}\overline{\kappa}^{+}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ $-\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$
$T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$ $-2(C_{S}+C_{DS})$
$3^{*}$ $T^{0}_{c\bar{s}}\to D^{0}\kappa^{0}$ $2C_{S}+C_{DS}$ $T^{0}_{c\bar{s}}\to D_{s}^{+}a_{0}^{-}$ $-(2C_{S}+C_{DS})$
$4^{*}$ $T^{0}_{cs}\to D^{0}\overline{\kappa}^{0}$ $2C_{S}+C_{DS}$ $T^{0}_{cs}\to D^{+}\kappa^{-}$ $-(2C_{S}+C_{DS})$
5 $T^{+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c}\to D^{+}a_{0}^{0}$ $-\frac{1}{2}(2C_{S}+C_{DS})$
$T^{+}_{c}\to D_{s}^{+}\overline{\kappa}^{0}$ $-\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c}\to D^{+}\sigma$ $\frac{1}{2}(\sqrt{2}c_{\phi}+s_{\phi})(2C_{S}+C_{DS})$
$T^{+}_{c}\to D^{+}f_{0}$ $-\frac{1}{2}(c_{\phi}-\sqrt{2}s_{\phi})(2C_{S}+C_{DS})$
6 $T^{0}_{c}\to D^{0}a_{0}^{0}$ $-\frac{1}{2}(2C_{S}+C_{DS})$ $T^{0}_{c}\to D^{+}a_{0}^{-}$ $-\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$
$T^{0}_{c}\to D_{s}^{+}\kappa^{-}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{0}_{c}\to D^{+}\sigma$ $-\frac{1}{2}(\sqrt{2}c_{\phi}+s_{\phi})(2C_{S}+C_{DS})$
@@
7 $T^{\prime\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{\sqrt{2}}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}a_{0}^{0}$ $-\frac{1}{2}(C_{DS}-2C_{S}+2ACC_{S})$
@@
$T^{\prime\prime+}_{c}\to D^{0}f_{0}$ $\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}\sigma$ $-\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$
$+\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S}$ $+\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$-ACC_{S}-3ACC_{D})$ $-ACC_{S}-3ACC_{D})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}f_{0}$ $\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime 0}_{c}\to D^{+}\sigma$ $-\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$
$+\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S}$ $+\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$-ACC_{S}-3ACC_{D})$ $-ACC_{S}-3ACC_{D})$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $-\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{3}(C_{DS}-2C_{S}+2ACC_{S})$
$+\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S}$ $+\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$-ACC_{S}-3ACC_{D})$ $-ACC_{S}-3ACC_{D})$
@@
$2^{\prime*}$ $T^{+}_{c\bar{s}}\to D^{0}\kappa^{+}$ $\frac{1}{\sqrt{2}}C_{DS}$ $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ $-\frac{1}{\sqrt{2}}C_{DS}$
@@
$11^{\prime*}$ $T^{0}_{c}\to D^{0}a_{0}^{0}$ $\frac{2}{\sqrt{3}}C_{DS}$ $T^{0}_{c}\to D^{+}a_{0}^{-}$ $-\sqrt{\frac{2}{3}}C_{DS}$
@@
$13^{\prime}$ $T^{\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{2\sqrt{3}}C_{DS}$ $T^{\prime+}_{c}\to D^{+}a_{0}^{0}$ $-\frac{1}{2\sqrt{6}}C_{DS}$
$T^{\prime+}_{c}\to D^{+}\sigma$ $-\frac{\sqrt{6}}{4}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime+}_{c}\to D^{+}f_{0}$ $-\frac{\sqrt{6}}{4}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$T^{\prime+}_{c}\to D_{s}^{+}\overline{\kappa}^{0}$ $-\frac{\sqrt{3}}{2}C_{DS}$
@@
$T^{\prime 0}_{c}\to D^{+}\sigma$ $-\frac{\sqrt{6}}{4}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime 0}_{c}\to D^{+}f_{0}$ $-\frac{\sqrt{6}}{4}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$T^{\prime 0}_{c}\to D_{s}^{+}\kappa^{-}$ $-\frac{\sqrt{3}}{2}C_{DS}$
@@
$T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $-\frac{1}{\sqrt{2}}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $-\frac{1}{\sqrt{2}}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$16^{\prime}$ $T^{\prime\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{2}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}a_{0}^{0}$ $-\frac{1}{2\sqrt{2}}(C_{DS}+4ACC_{S})$
@@
$T^{\prime\prime+}_{c}\to D^{+}f_{0}$ $\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{12}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}\sigma$ $-\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{12}(C_{DS}+4ACC_{S})$
$-\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $-\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}f_{0}$ $\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{12}(C_{DS}+4ACC_{S})$ $T^{\prime\prime 0}_{c}\to D^{+}\sigma$ $-\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{12}(C_{DS}+4ACC_{S})$
$-\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $-\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $-\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{6}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{6}(C_{DS}+4ACC_{S})$
$-\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $-\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
The sizes of the amplitudes of the following decay modes of flavor exotic states, $T^{++}_{c\bar{s}}\to D^{+}\kappa^{+},D_{s}^{+}a_{0}^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$, $T^{0}_{c\bar{s}}\to D^{0}\kappa^{0},D_{s}^{+}a_{0}^{-}$ and $T^{0}_{cs}\to D^{0}\overline{\kappa}^{0},D^{+}\kappa^{-}$, are identical, giving $|A|=|2C_{S}+C_{DS}|$, and factor $1/\sqrt{2}$ are needed for $|A|$ of $T^{+}_{c\bar{s}}\to D^{0}\overline{\kappa}^{+}$ and $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ decays. The flavor exotic nature of most of these states is self-evident from the table. For example, the $T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$ decay reveals the exotic quantum number of $T^{+}_{c\bar{s}}$.
@@ IV.1.1 Topological amplitudes in $\overline{B}\to D\overline{T}$ decays
The $\overline{B}\to D\overline{T}$ decays in $\Delta S=-1$ transitions are governed by the $(\bar{s}c)_{V-A}(\bar{c}b)_{V-A}$ operator. The flavor structure of the above operator can be expressed as $(\bar{q}_{j}H^{j}c)\,(\bar{c}b)$ with $\bar{q}_{j}=(\bar{u},\,\bar{d},\,\bar{s})$ and
@@
Consequently, in scenario I the Hamiltonian governing $\overline{B}\to D\overline{T}(\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}])$ decays is given by
$H_{\rm eff}$ $=$ $T_{\bar{T}}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{[ji]}+C_{\bar{T}}\,\overline{B}_{m}T^{m}_{[jk]}H^{j}\overline{D}^{k}+E_{\bar{T}\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{[ik]}\overline{D}^{k},$ (34)
@@
$H_{\rm eff}$ $=$ $T_{\bar{T}}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{\{ji\}}+C_{\bar{T}}\,\overline{B}_{m}T^{m}_{\{jk\}}H^{j}\overline{D}^{k}+E_{\bar{T}\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{\{ik\}}\overline{D}^{k}.$ (35)
For notational simplicity, we use the same notations on topological amplitudes in both scenarios, without any assumption on their equality. There are three topological amplitudes, $T_{\bar{T}}$, $C_{\bar{T}}$ and $E_{\bar{T}\bar{B}}$, as depicted in Fig. 4. These $T_{\bar{T}}$, $C_{\bar{T}}$ and $E_{\bar{T}\bar{B}}$ amplitudes denote external $W$, internal $W$ and exchange amplitudes, respectively, where the subscripts denote the particle(s) receiving the $\bar{c}s$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}s$ transition.
Figure 4: Topological diagrams of (a) $C_{\bar{T}}$ (internal $W$), (b) $T_{\bar{T}}$ (external $W$)and (c) $E_{\bar{T}\bar{B}}$ (exchange), amplitudes in $\overline{B}\to D\overline{T}$ decays. The subscripts indicate the final state meson(s) receiving the $\bar{c}q_{j}$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}q_{j}$ transition. These are flavor flow diagrams. Only (a) contributes to modes with flavor exotic states or other states in the same multiplets.
@@
$|T_{\bar{T}}|>|C_{\bar{T}}|>|E_{\bar{T}\bar{B}}|.$ (36)
Although the external $W$ diagram is expected to dominate, it is unable to produce a flavor exotic $\overline{T}$ in the final state, as the $q\bar{q}$ pair is flavor connected, see Fig. 4 (b). These flavor exotic $\overline{T}$ and other states in the same multiplets, namely ${\bf 6}$ in scenario I and $\overline{\bf 15}$ in scenario II, only have contribution from topological amplitude $C_{\bar{T}}$, and, consequently, their $\overline{B}\to D\overline{T}$ decay amplitudes are highly related.

### v3

@@ III.2.1 Topological amplitudes in $T\to DS$ decays
the mixing angle is taken to be $\phi=174.6^{\circ}$Maiani:2004uc . For simplicity of notation, we do not distinguish the topological amplitudes in scenario I and scenario II, which, however, do not imply that they are identical. We follow refs. Jaffe:1976ig ; Jaffe:1976ih on the quark contents of these low-lying scalar particles. For more information on the flavor structure of these particles, see Appendix C.
@@ III.2.2 $T\to DS$ decay amplitudes
$1^{*}$ $T^{++}_{c\bar{s}}\to D^{+}\kappa^{+}$ $2C_{S}+C_{DS}$ $T^{++}_{c\bar{s}}\to D_{s}^{+}a_{0}^{+}$ $(2C_{S}+C_{DS})$
$2^{*}$ $T^{+}_{c\bar{s}}\to D^{0}\overline{\kappa}^{+}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$
$T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$ $2(C_{S}+C_{DS})$
$3^{*}$ $T^{0}_{c\bar{s}}\to D^{0}\kappa^{0}$ $2C_{S}+C_{DS}$ $T^{0}_{c\bar{s}}\to D_{s}^{+}a_{0}^{-}$ $(2C_{S}+C_{DS})$
$4^{*}$ $T^{0}_{cs}\to D^{0}\overline{\kappa}^{0}$ $2C_{S}+C_{DS}$ $T^{0}_{cs}\to D^{+}\kappa^{-}$ $(2C_{S}+C_{DS})$
5 $T^{+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c}\to D^{+}a_{0}^{0}$ $\frac{1}{2}(2C_{S}+C_{DS})$
$T^{+}_{c}\to D_{s}^{+}\overline{\kappa}^{0}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{+}_{c}\to D^{+}\sigma$ $\frac{1}{2}(\sqrt{2}c_{\phi}+s_{\phi})(2C_{S}+C_{DS})$
$T^{+}_{c}\to D^{+}f_{0}$ $\frac{1}{2}(c_{\phi}-\sqrt{2}s_{\phi})(2C_{S}+C_{DS})$
6 $T^{0}_{c}\to D^{0}a_{0}^{0}$ $\frac{1}{2}(2C_{S}+C_{DS})$ $T^{0}_{c}\to D^{+}a_{0}^{-}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$
$T^{0}_{c}\to D_{s}^{+}\kappa^{-}$ $\frac{1}{\sqrt{2}}(2C_{S}+C_{DS})$ $T^{0}_{c}\to D^{+}\sigma$ $\frac{1}{2}(\sqrt{2}c_{\phi}+s_{\phi})(2C_{S}+C_{DS})$
@@
7 $T^{\prime\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{\sqrt{2}}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}a_{0}^{0}$ $\frac{1}{2}(C_{DS}-2C_{S}+2ACC_{S})$
@@
$T^{\prime\prime+}_{c}\to D^{0}f_{0}$ $\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}\sigma$ $\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$
$\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S} +$$\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$ACC_{S}-3ACC_{D})$ $ACC_{S}-3ACC_{D})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}f_{0}$ $\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime 0}_{c}\to D^{+}\sigma$ $\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{6}(C_{DS}-2C_{S}+2ACC_{S})$
$\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S} +$$\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$ACC_{S}-3ACC_{D})$ $ACC_{S}-3ACC_{D})$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $\frac{(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}-2C_{S}+2ACC_{S})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $\frac{(s_{\phi}+\sqrt{2}c_{\phi})}{3}(C_{DS}-2C_{S}+2ACC_{S})$
$\frac{2(\sqrt{2}s_{\phi}+2c_{\phi})}{3}(C_{DS}+C_{S} +$$\frac{2(\sqrt{2}c_{\phi}-2s_{\phi})}{3}(C_{DS}+C_{S}$
$ACC_{S}-3ACC_{D})$ $ACC_{S}-3ACC_{D})$
@@
$2^{\prime*}$ $T^{+}_{c\bar{s}}\to D^{0}\kappa^{+}$ $\frac{1}{\sqrt{2}}C_{DS}$ $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ $\frac{1}{\sqrt{2}}C_{DS}$
@@
$11^{\prime*}$ $T^{0}_{c}\to D^{0}a_{0}^{0}$ $\frac{2}{\sqrt{3}}C_{DS}$ $T^{0}_{c}\to D^{+}a_{0}^{-}$ $\sqrt{\frac{2}{3}}C_{DS}$
@@
$13^{\prime}$ $T^{\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{2\sqrt{3}}C_{DS}$ $T^{\prime+}_{c}\to D^{+}a_{0}^{0}$ $\frac{1}{2\sqrt{6}}C_{DS}$
$T^{\prime+}_{c}\to D^{+}\sigma$ $\frac{\sqrt{6}}{4}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime+}_{c}\to D^{+}f_{0}$ $\frac{\sqrt{6}}{4}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$T^{\prime+}_{c}\to D_{s}^{+}\overline{\kappa}^{0}$ $\frac{\sqrt{3}}{2}C_{DS}$
@@
$T^{\prime 0}_{c}\to D^{+}\sigma$ $\frac{\sqrt{6}}{4}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime 0}_{c}\to D^{+}f_{0}$ $\frac{\sqrt{6}}{4}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$T^{\prime 0}_{c}\to D_{s}^{+}\kappa^{-}$ $\frac{\sqrt{3}}{2}C_{DS}$
@@
$T^{\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $\frac{1}{\sqrt{2}}(\sqrt{2}c_{\phi}+s_{\phi})C_{DS}$ $T^{\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $\frac{1}{\sqrt{2}}(\sqrt{2}s_{\phi}-c_{\phi})C_{DS}$
$16^{\prime}$ $T^{\prime\prime+}_{c}\to D^{0}a_{0}^{+}$ $\frac{1}{2}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}a_{0}^{0}$ $\frac{1}{2\sqrt{2}}(C_{DS}+4ACC_{S})$
@@
$T^{\prime\prime+}_{c}\to D^{+}f_{0}$ $\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{12}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c}\to D^{+}\sigma$ $\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{12}(C_{DS}+4ACC_{S})$
$\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
@@
$T^{\prime\prime 0}_{c}\to D^{0}f_{0}$ $\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{12}(C_{DS}+4ACC_{S})$ $T^{\prime\prime 0}_{c}\to D^{+}\sigma$ $\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{12}(C_{DS}+4ACC_{S})$
$\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
@@
$T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}f_{0}$ $\frac{\sqrt{2}c_{\phi}-2s_{\phi}}{6}(C_{DS}+4ACC_{S})$ $T^{\prime\prime+}_{c\bar{s}}\to D_{s}^{+}\sigma$ $\frac{\sqrt{2}s_{\phi}+2c_{\phi}}{6}(C_{DS}+4ACC_{S})$
$\frac{4(\sqrt{2}c_{\phi}+s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$ $\frac{4(c_{\phi}-\sqrt{2}s_{\phi})}{3}(C_{DS}+ACC_{S}+3ACC_{D})$
The sizes of the amplitudes of the following decay modes of flavor exotic states, $T^{++}_{c\bar{s}}\to D^{+}\kappa^{+},D_{s}^{+}a_{0}^{+}$, $T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$, $T^{0}_{c\bar{s}}\to D^{0}\kappa^{0},D_{s}^{+}a_{0}^{-}$ and $T^{0}_{cs}\to D^{0}\overline{\kappa}^{0},D^{+}\kappa^{-}$, are identical, giving $|A|=|2C_{S}+C_{DS}|$, and factor $1/\sqrt{2}$ are needed for $|A|$ of $T^{+}_{c\bar{s}}\to D^{0}\overline{\kappa}^{+}$ and $T^{+}_{c\bar{s}}\to D^{+}\kappa^{0}$ decays. The flavor exotic nature of most of these states is self-evident from their decay modes shown in the table. For example, the $T^{+}_{c\bar{s}}\to D_{s}^{+}a_{0}^{0}$ decay reveals the exotic quantum number of $T^{+}_{c\bar{s}}$, as $a_{0}$ is an iso-vector.
@@ IV.1.1 Topological amplitudes in $\overline{B}\to D\overline{T}$ decays
The $\overline{B}\to D\overline{T}$ decays in $\Delta S=-1$ transitions are governed by the a tree operator $(\bar{s}c)_{V-A}(\bar{c}b)_{V-A}$ and a penguin operator $(\bar{c}c)(\bar{s}b)$. These operators have the same flavor structure that can be expressed as $(\bar{q}_{j}H^{j}c)\,(\bar{c}b)$ with $\bar{q}_{j}=(\bar{u},\,\bar{d},\,\bar{s})$ and
@@
Consequently, in scenario I the Hamiltonian governing the $\overline{B}\to D\overline{T}(\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}])$ decays is given by
$H_{\rm eff}$ $=$ $T_{\bar{T}}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{[ji]}+C_{\bar{T}}\,\overline{B}_{m}T^{m}_{[jk]}H^{j}\overline{D}^{k}+E_{\bar{T}\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{[ik]}\overline{D}^{k}$ (50)
$+P_{\bar{T},1}\,\overline{B}_{m}T^{m}_{[jk]}H^{j}\overline{D}^{k}+P_{\bar{T},2}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{[ji]}+PA_{\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{[ik]}\overline{D}^{k},$
@@
$H_{\rm eff}$ $=$ $T_{\bar{T}}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{\{ji\}}+C_{\bar{T}}\,\overline{B}_{m}T^{m}_{\{jk\}}H^{j}\overline{D}^{k}+E_{\bar{T}\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{\{ik\}}\overline{D}^{k}$ (51)
$+P_{\bar{T},1}\,\overline{B}_{m}T^{m}_{\{jk\}}H^{j}\overline{D}^{k}+P_{\bar{T},2}\,\overline{B}_{m}\overline{D}^{m}H^{j}T^{i}_{\{ji\}}+PA_{\bar{B}}\,\overline{B}_{j}H^{j}T^{i}_{\{ik\}}\overline{D}^{k}.$
For notational simplicity, we use the same notations on topological amplitudes in both scenarios, without any assumption on their equality. There are three tree topological amplitudes, $T_{\bar{T}}$, $C_{\bar{T}}$ and $E_{\bar{T}\bar{B}}$ and three penguin topological amplitudes, $P_{\bar{T},1}$, $P_{\bar{T},2}$ and $PA_{\bar{B}}$, as depicted in Fig. 4. These $T_{\bar{T}}$, $C_{\bar{T}}$ and $E_{\bar{T}\bar{B}}$ amplitudes denote external $W$, internal $W$ and exchange amplitudes, respectively, where the subscripts denote the particles receiving the $\bar{c}s$ quarks from the $W$ line in the $b\to cW^{-},W^{-}\to\bar{c}s$ transition. For $P_{\bar{T},1}$, $P_{\bar{T},2}$ and $PA_{\bar{B}}$ amplitudes, they are penguin and penguin-annihilation amplitudes, respectively, where the subscripts denote the particles receiving the $s$ quark from the $b\to s\bar{c}c$ transition.
Figure 4: Topological diagrams of (a) $C_{\bar{T}}$ (internal $W$), (b) $T_{\bar{T}}$ (external $W$), (c) $E_{\bar{T}\bar{B}}$ (exchange), (d) $P_{\bar{T},1}$ (penguin), (e) $P_{\bar{T},2}$ (penguin)and (f) $PA_{\bar{B}}$ (penguin annihilation) amplitudes in $\overline{B}\to D\overline{T}$ decays. These are flavor flow diagrams. Only (a) and (d) contribute to modes with flavor exotic states or other states in the same multiplets.
@@
$|T_{\bar{T}}|>|C_{\bar{T}}|>|P_{\bar{T},1}|,|P_{\bar{T},2}|,|E_{\bar{T}\bar{B}}|,|PA_{\bar{B}}|.$ (52)
Although the external $W$ diagram is expected to dominate, it is unable to produce a flavor exotic $\overline{T}$ in the final state, as the $q\bar{q}$ pair is flavor connected, see Fig. 4 (b). These flavor exotic $\overline{T}$ and other states in the same multiplets, namely $\bf 6$} in scenario I and $\overline{\bf 15}$ in scenario II, only have contribution from $(C_{\bar{T}}+P_{\bar{T},1})$, [^1] These two tpological amplitudes can only occur in this combination, see Eqs. (50) and (51). and, consequently, their $\overline{B}\to D\overline{T}$ decay amplitudes are highly related.